OpenStudy (anonymous):

Find two numbers whose sum is 40, such that the sum of the square of one number plus 6 times the other number is a minimum.

OpenStudy (raden):

let that number are a and b respectively. according information above, we get the equation : x+y = 40 asking the minimum value of K (assumed) = x^2 + 6y from the first equation x+y = 40 ----> y = 40 - x. subtitute this into K, get : K = x^2 + 6y K = x^2 + 6(40-x) K = x^2 - 6x + 240 K has the minimum value when x = -b/2a a = coefficient of x^2 while b = coefficient of x. so, x = -b/2a = -(-6)/(2*1) = 3 because y = 40 - x, subtitute x = 3 into y we get y = 40 - 3 = 37 thus, they are x = 3 and y =37